f(x)= x^2, g(x)= 3x + 1
find (f º g)(2)

(gº f)(2)

Answer by robertb(5830)

g(2) = 3*2+1 = 7 ==> f o g(2) = f(2) = 49
f(2) = 4 ==> g o f (2) = g(4) = 3*4 + 1 = 13